# Baryon Pasting the Uchuu Lightcone Simulation

### v1

@@ 5 Discussion
Our findings are subject to several limitations that will be investigated further in future work. Firstly, the BP feedback model encapsulates the SNe and AGN feedback into one unified parameter. This approach does not adequately capture the intricate interactions occurring between SNe and AGN feedback (e.g., Medlock et al., 2024). Leveraging hydrodynamical cosmological simulations that incorporate varied SNe and AGN feedback physics across a broad spectrum of mass scales, such as those from CAMELS (Villaescusa-Navarro et al., 2021) and CarpoolGP (Lee et al., 2024), could enhance the feedback models applied in our study.
@@ 6 Conclusions
7 Acknowledgements
This work is supported by NASA ATP23-0154 grant and the Yale Center for Research Computing. A.F. acknowledges support from the National Science Foundation under Cooperative Agreement AST-2421782. M.S. acknowledges support from MEXT KAKENHI Grant Number (20H05861, 23K19070, 24H00215, 24H00221). KO is supported by JSPS KAKENHI Grant Number JP22K14036 and JP24H00215. T.I. has been supported by IAAR Research Support Program in Chiba University Japan, MEXT/JSPS KAKENHI (Grant Number JP19KK0344 and JP23H04002), MEXT as “Program for Promoting Researches on the Supercomputer Fugaku” (JPMXP1020230406), and JICFuS. HM is supported by JSPS KAKENHI Grand Numbers JP20H01932, JP23H00108, and 22K21349, and Tokai Pathways to Global Excellence (T-GEx), part of MEXT Strategic Professional Development Program for Young Researchers We thank Instituto de Astrofisica de Andalucia (IAA-CSIC), Centro de Supercomputacion de Galicia (CESGA) and the Spanish academic and research network (RedIRIS) in Spain for hosting Uchuu DR1, DR2 and DR3 in the Skies & Universes site for cosmological simulations. The Uchuu simulations were carried out on Aterui II supercomputer at Center for Computational Astrophysics, CfCA, of National Astronomical Observatory of Japan, and the K computer at the RIKEN Advanced Institute for Computational Science. The Uchuu Data Releases efforts have made use of the skun@IAA_RedIRIS and skun6@IAA computer facilities managed by the IAA-CSIC in Spain (MICINN EU-Feder grant EQC2018-004366-P).
@@ A.1 Profile Comparison
In the bottom panels of the same figure, we compare the Universal Pressure Profile between the Planck-XMM measurements (Planck Collaboration et al., 2013) and the BP model for the same halos. Again there is good agreement between the profiles of the BP model and the Universal Pressure Profile. Note that we do not fit the BP model to the Planck-XMM data. All the model parameters are from our calibration with the Chandra-SPT data, except the inner Polytropic index where we change from the fiducial value of $0.10$ to $1.0$. Note that the thermal pressure profiles are normalized by $P_{500c}=1.45\times 10^{-11}\,{\rm erg\,cm^{-3}}\left(\frac{f_{b}}{0.174}\right)\left(\frac{h}{0.7}\right)^{2}E(z)^{8/3}\left(\frac{M_{500c}}{10^{15}h^{-1}M_{\odot}}\right)^{2/3}$ to account for self-similar mass and redshift dependence in the normalization of the pressure profile.
@@
In Figure 12 we compare the Halo model-based tSZ power spectra computed with the BP code to the tSZ power spectrum from Planck (Planck Collaboration et al., 2016), ACT, and SPT. The BP tSZ power spectrum is computed with mass range $\log_{10}(M_{\rm vir}/M_{\odot})=[13,16]$ and redshift range $z=[0.0,2.0]$, following Equations (39)-(42) in Osato & Nagai (2023), with varying non-thermal pressure fraction parameter $A_{\rm nt}$. We use the Planck 2018 cosmology (Planck Collaboration et al., 2020) to compute the power spectrum. We compare the model power spectra with that of Planck from Planck Collaboration et al. (2016). The BP tSZ power spectrum with the fiducial non-thermal pressure fraction parameter $A_{\rm nt}=0.45$ calibrated from cosmological hydrodynamical simulation provides a good match to the Planck measurements. This is also consistent with previous works that compared the BP model power spectra with observations (Osato et al., 2018).

### v2

@@ 5 Discussion
Our findings are subject to several limitations that will be investigated further in future work. Firstly, the BP feedback model encapsulates the SNe and AGN feedback into one unified parameter. This approach does not adequately capture the intricate interactions that occur between SNe and AGN feedback (e.g., Medlock et al., 2024). Leveraging hydrodynamical cosmological simulations that incorporate varied SNe and AGN feedback physics across a broad spectrum of mass scales, such as those from CAMELS (Villaescusa-Navarro et al., 2021) and CarpoolGP (Lee et al., 2024), could enhance the feedback models applied in our study.
@@ 6 Conclusions
Acknowledgements
We thank the anonymous referee for their comments and feedback. This work is supported by NASA ATP23-0154 grant and the Yale Center for Research Computing. A.F. acknowledges support from the National Science Foundation under Cooperative Agreement AST-2421782 the Simons Foundation award MPS-AI-00010515. M.S. acknowledges support from MEXT KAKENHI Grant Number (20H05861, 23K19070, 24H00215, 24H00221). KO is supported by JSPS KAKENHI Grant Number JP22K14036 and JP24H00215. T.I. has been supported by IAAR Research Support Program in Chiba University Japan, MEXT/JSPS KAKENHI (Grant Number JP19KK0344 and JP23H04002), MEXT as “Program for Promoting Researches on the Supercomputer Fugaku” (JPMXP1020230406), and JICFuS. HM is supported by JSPS KAKENHI Grand Numbers JP20H01932, JP23H00108, and 22K21349, and Tokai Pathways to Global Excellence (T-GEx), part of MEXT Strategic Professional Development Program for Young Researchers We thank Instituto de Astrofisica de Andalucia (IAA-CSIC), Centro de Supercomputacion de Galicia (CESGA) and the Spanish academic and research network (RedIRIS) in Spain for hosting Uchuu DR1, DR2 and DR3 in the Skies & Universes site for cosmological simulations. The Uchuu simulations were carried out on Aterui II supercomputer at Center for Computational Astrophysics, CfCA, of National Astronomical Observatory of Japan, and the K computer at the RIKEN Advanced Institute for Computational Science. The Uchuu Data Releases efforts have made use of the skun@IAA_RedIRIS and skun6@IAA computer facilities managed by the IAA-CSIC in Spain (MICINN EU-Feder grant EQC2018-004366-P).
@@ A.1 Profile Comparison
In the bottom panels of the same figure, we compare the pressure profile between the fit to observations from Planck and SPT-SZ (Melin & Pratt 2023, see also He et al. 2021) for another updated model fit to the Planck measurements Planck Collaboration et al. 2013), and the BP model for the same halos. Again there is good agreement between the profiles of the BP model and the Universal Pressure Profile. Note that we do not fit the BP model to the Planck-XMM data. All the model parameters are from our calibration with the Chandra-SPT data, except the inner Polytropic index where we change from the fiducial value of $0.10$ to $1.0$. Note that the thermal pressure profiles are normalized by $P_{500c}=1.45\times 10^{-11}\,{\rm erg\,cm^{-3}}\left(\frac{f_{b}}{0.174}\right)\left(\frac{h}{0.7}\right)^{2}E(z)^{8/3}\left(\frac{M_{500c}}{10^{15}h^{-1}M_{\odot}}\right)^{2/3}$ to account for self-similar mass and redshift dependence in the normalization of the pressure profile.
@@
In Figure 12 we compare the Halo model-based tSZ power spectra computed with the BP code to the tSZ power spectrum from Planck (Planck Collaboration et al., 2016), ACT, and SPT. The BP tSZ power spectrum is computed with mass range $\log_{10}(M_{\rm vir}/M_{\odot})=[13,16]$ and redshift range $z=[0.0,2.0]$, following Equations (39)-(42) in Osato & Nagai (2023), with varying non-thermal pressure fraction parameter $A_{\rm nt}$. We use the Planck 2018 cosmology (Planck Collaboration et al., 2020) to compute the power spectrum. We compare the model power spectra with that of Planck (Planck Collaboration et al., 2016), ACT (Dunkley et al., 2013), and SPT (George et al., 2015). The BP tSZ power spectrum with the fiducial non-thermal pressure fraction parameter $A_{\rm nt}=0.45$ calibrated from cosmological hydrodynamical simulation provides a good match to the Planck measurements. This is also consistent with previous works that compared the BP model power spectra with observations (Osato et al., 2018).